OpenStudy (mallorysipp234):

If cos Θ = 4/7, what are the values of sin Θ and tan Θ?

OpenStudy (mallorysipp234):

@mathmale @satellite73

OpenStudy (mallorysipp234):

@saifoo.khan

OpenStudy (mallorysipp234):

@ranga

OpenStudy (mallorysipp234):

@ShaTay

OpenStudy (anonymous):

I have No clue... sorry

OpenStudy (ranga):

Use the identity sin^2(x) + cos^2(x) = 1. Plug in the value of cos(x) into the above identity and solve for sin(x). Them, tan(x) = sin(x) / cos(x).

OpenStudy (anonymous):

can't you use a calculator? in scientific mode

OpenStudy (mallorysipp234):

I still dont understand.

OpenStudy (ranga):

sin^2(x) + cos^2(x) = 1 cos(x) = 4/7 sin^2(x) + (4/7)^2 = 1 sin^2(x) + 16/49 = 1 sin^2(x) = 1 - 16/49 = (49-16)/49 = 33/49 sin(x) = sqrt(33/49) = sqrt(33) / 7. tan(x) = sin(x) / cos(x) = sqrt(33) / 7 * 7 / 4 = sqrt(33) / 44

OpenStudy (mallorysipp234):

@ranga

OpenStudy (ranga):

You can also solve this problem by drawing a triangle, marking one of the non-ninety degree angle as x, using the definition of cos(x) to identify the adjacent and hypotenuse, etc.: 

OpenStudy (ranga):

In my reply before last, tan(x) = sqrt(33)/4 (not 44).